OpenStudy (anonymous):

We have a bag with 6 red marbles, 1 blue marbles, and 2 green marbles in it. What is the probability of randomly selecting a Red marble from the bag?

OpenStudy (a_clan):

6/9 =2/3

OpenStudy (anonymous):

correct

OpenStudy (tad1):

There are 9 marbles in the bag. There are 6 red marbles. so the probability of drawing a red marble is 6/9=2/3